OpenStudy (anonymous):

A string is wound around the outside of a uniform disc of mass M and radius R. If the disc is released from rest, calculate the acceleration of the disc as the string unwinds from the disc in terms of Ag.

OpenStudy (anonymous):



OpenStudy (vincent-lyon.fr):

What method have you chosen to solve this? (forces, energy?)

OpenStudy (anonymous):

 equation of motions are \[Ma = Mg - T ~~~~(1)\] \[\frac{1}{2}MR^2 \frac{d \omega}{dt}=RT~~~~(2)\] \[a=R \frac{d \omega }{dt}~~~~~~(3)\] (1)+(2)/R and (3) We get \[\frac{3}{2}Ma=Mg\] therefore \[a=\frac{2}{3}g\]